# An All-digital 65-nm Tsetlin Machine Image Classification Accelerator with 8.6 nJ per MNIST Frame at 60.3k Frames per Second

### v1

@@ Abstract
We present an all-digital programmable machine learning accelerator chip for image classification, underpinning on the Tsetlin machine (TM) principles. The TM is a machine learning algorithm founded on propositional logic, utilizing sub-pattern recognition expressions called clauses. The accelerator implements the coalesced TM version with convolution, and classifies booleanized images of 28$\times$28 pixels with 10 categories. A configuration with 128 clauses is used in a highly parallel architecture. Fast clause evaluation is obtained by keeping all clause weights and Tsetlin automata (TA) action signals in registers. The chip is implemented in a 65 nm low-leakage CMOS technology, and occupies an active area of 2.7 mm2. At a clock frequency of 27.8 MHz, the accelerator achieves 60.3k classifications per second, and consumes 8.6 nJ per classification. The latency for classifying a single image is 25.4 $\mu$s which includes system timing overhead. The accelerator achieves 97.42%, 84.54% and 82.55% test accuracies for the datasets MNIST, Fashion-MNIST and Kuzushiji-MNIST, respectively, matching the TM software models.
@@ I Introduction
Hardware (HW) accelerators are commonly integrated in modern electronic systems to relieve the burden on system processors from extensive application-specific workloads. Performance and energy efficiency within a smaller form factor are key considerations for accelerator designs, particularly for Internet-of-Things (IoT), where the compute and memory resources of edge nodes typically are limited. In addition, such nodes are often battery-operated, and it is critical to minimize the overall energy footprints for longer battery operational lifetime.
@@
Current methods of accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3]. CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources. As such, they require large number of micro-architectural operations for each classification and have significant energy footprints [3]. The complexity and energy consumption can be reduced by compromising the data precision in CNNs [4]. Binary and ternary activations and weights have been applied in several integrated circuit (IC) solutions that have shown good test accuracy with very low power consumption [5, 6, 7, 8, 9].
Vision Transformer (ViT) [10] solutions obtain the highest score on image classification benchmarks, but they are computationally more intensive than CNNs [11]. ViT-based neural networks utilize a global receptive field that captures correlations between patches in an image [10, 11]. This is in contrast to CNNs, where convolutional layers extract features of an image in a hierarchical manner, with a limited local receptive field. Lightweight Transformer models have been developed, e.g., MobileViT [12], that have been applied for HW implementation [13].
@@
Unlike the numerical models based on neural networks mentioned above, the Tsetlin machine [14] is an ML algorithm founded on propositional logic. During training, parallel conjunctive clauses are defined by independent learning automata. Each clause represents a logical sub-pattern and will output a logical 1 if it matches an input sample, and 0 otherwise. The TM features a single-layer structure with highly interpretable outputs [15, 16, 17], making it well-suited for application areas such as biomedical, cybernetics, IoT, and law, where interpretability is essential. Furthermore, because its primary operations are Boolean, TM solutions are both low-complexity and hardware-efficient. The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training. This makes TMs ideal for low-energy application-specific integrated circuit (ASIC) implementations [18]. Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [19, 20].
When compared against deep neural networks (DNNs) and other ML algorithms, the TM has demonstrated competitive performance. It has been tested on tabular data [14], natural language [17, 15], images [21, 22, 23] and regression [24]. Through the works in [25, 26, 27] the convergence of the TM’s basic operators has been verified.
@@
The original TM version [14] is called the vanilla TM. For image classification, improved test accuracy is achieved by the convolutional TM (CTM) [21]. The CTM has obtained a peak test accuracy of 99.4% on the MNIST dataset [28]. For the Fashion-MNIST (FMNIST) [29] and Kuzushiji-MNIST (KMNIST) [30] datasets, the peak test accuracies achieved are 91.5% and 96.31% respectively. Ensemble TM approaches have shown promising accuracies for more complex image datasets, such as CIFAR-10 and CIFAR-100, which are currently being investigated. In the case of CIFAR-10, the approach denoted TM Composites [31] has demonstrated a test accuracy of 82.8% [32].
@@
In this paper, we present an all-digital inference accelerator ASIC based on the coalesced TM (CoTM) [22]. The CoTM utilizes a single clause pool, in contrast to the vanilla TM which uses one pool of clauses per class. As we operate the CoTM with convolution, we here denote it as a convolutional CoTM (ConvCoTM). For configurations with few clauses, the ConvCoTM achieves better accuracy than the vanilla CTM, which makes the ConvCoTM especially attractive for HW solutions with constrained resources.
The main motivation behind the development of this ASIC was to demonstrate the HW friendliness and power efficiency of a ConvCoTM-based accelerator. The main contributions of this paper are:
@@
  • The ConvCoTM accelerator core operates with an energy per classification (EPC) of 8.6 nJ. To the best of the authors’ knowledge, this is the lowest reported EPC for manufactured ASICs that operate on a similar test dataset and with comparable test accuracy.
@@
  • The ConvCoTM solution is a fully digital, synchronous design compatible with standard digital workflows.
@@ II Related Work
An architecture for a binary-weight CNN is described in [8]. Images of size $224\times 224\times 3$ are classified at a rate of 11 frames per second (FPS). The technology used is 65 nm CMOS and simulations show a power consumption of 895 $\mu$W at 0.6 V.
The accelerator ASIC in [35] is manufactured in 65 nm CMOS. Its classification rates for Imagenet [36] are 35/0.7 FPS for the AlexNet and VGG-16 models respectively. The ASIC consumes 278 mW/236 mW with corresponding EPC values of 7.9 mW/0.4 mJ. The successor to [35] is the work described in [37]. Here, additional steps are taken to reduce energy consumption further. In particular, it reduces the movement of all data types during the various operational steps of the implemented CNN algorithm. A highly flexible on-chip network is utilized and sparse data are processed efficiently. When operating on $224\times 224\times 3$ images from Imagenet, simulations show a power consumption of 1.44 mW at 1470.6 FPS. The technology used is 65 nm CMOS.
@@
In-memory-computing (IMC) solutions can provide very high energy efficiency, as the bottleneck related to memory access is relieved. IMC has especially been utilized for analog/mixed-signal ML accelerator designs. An IMC mixed-signal 28 nm CMOS BNN processor is described in [5]. This accelerator achieves an EPC of 3.8 $\mu$J when operating on CIFAR-10. The classification rate is 237 FPS with a power consumption of 0.9 mW.
@@
In [9] a charge-domain IMC ternary neural network (TNN) classifier is described. The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model. The accelerator classifies MNIST images at a rate of 549 FPS. It has very low power consumption, only 96 $\mu$W, which corresponds to 0.18 $\mu$J per classification. The test accuracy achieved is 97.1%.
@@
Neuromorphic approaches, utilizing spiking neural networks (SNNs) have achieved very low energy consumption. In [38] a mixed-signal neuromorphic SNN IC is reported, which has been manufactured in a 65 nm CMOS technology. It achieves a test accuracy on MNIST of 95.35%, a power consumption of 0.517 mW and an EPC of 12.92 nJ. Simulation results of a 28 nm CMOS charge-domain computing SNN solution are reported in [39]. The accelerator achieves an EPC of 15.09 nJ, a latency of 0.46 $\mu$s and a test accuracy on MNIST of 94.81%. In [40] simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described. On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC.
@@
For the TM-based hardware solutions, the first reported IC is described in [18]. It utilizes the vanilla TM architecture and operates on a 3-class machine learning (ML) problem. The IC is implemented in a 65 nm CMOS technology and supports both training and inference. In [19], the first TM-based HW accelerator with convolution is reported. It is an FPGA solution which operates on a 2-class pattern recognition problem (two-dimensional noisy XOR [21]) in 4$\times$4 Boolean images, and employs a 2$\times$2 convolution window. Full on-device training is implemented.
The FPGA solution reported in [41] is based on the vanilla TM and stores and processes only the include signals of the TA actions, utilizing the typical TM model’s high sparsity. It therefore represents a highly compact implementation when it comes to HW resources. The design operates on the MNIST dataset, with sequential processing of each clause and each literal.
@@
In [20] the first reported ConvCoTM-based HW solution is presented. It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes complete on-device training. For the MNIST dataset a test accuracy of 97.6% is achieved, with a classification rate of 134k samples per second, and an EPC of 13.3 $\mu$J.
@@
Several articles describe the design and simulation of HW solutions for TM accelerators. In [42, 43, 44] low-latency asynchronous TM solutions for inference and training are presented. A mixed-signal IMC TM solution is described in [45] with an EPC for MNIST of 13.9nJ. It is an inference solution with a Boolean-to-current architecture, and resistive RAM (ReRAM) transistor cells are employed to eliminate the need for energy-hungry analog-to-digital and digital-to-analog conversions. Another IMC TM concept based on Y-flash cells is reported in [46].
Based on the above literature review, developing a ConvCoTM ASIC solution is a critical step toward achieving higher energy efficiency, performance, and scalability for TM-based hardware accelerators. Existing FPGA implementations, while effective, are limited in energy efficiency. Comparatively, in-memory computing (IMC) and binary neural network (BNN)-based ASICs demonstrate significantly lower energy consumption and higher operational efficiency, underscoring the potential benefits of an ASIC approach. By leveraging a standard digital design flow and CMOS technology, an ASIC solution for ConvCoTM could deliver real-time, power-efficient classification. This would position ConvCoTM as a competitive alternative to CNNs and BNNs in resource-constrained environments.
@@ III-A General TM Concepts
The Tsetlin Automaton (TA) of two-action type [47] constitutes the basic learning element in a TM. The TA will either include or exclude a literal in a conjunctive clause, depending on its state. One team of TAs is applied per clause, and the different TA actions are obtained during the training process which involves feedback mechanisms to each TA [14]. For an inference-only solution, as reported in this paper, we need only the TA action signals from the trained model, not the complete TAs.
In a TM the number of clauses, $n$, is a user specified integer. The output of a single clause, $c_{j}$, where $j\in\{0,\ldots,n-1\}$, is given by
@@
where $l_{k}$ is the literal with index k. k belongs to $I_{j}\subseteq\{0,....,2o-1\}$, where $I_{j}$ is the set of literal indexes for clause $j$ where the TA action is include [14].
@@ III-B CoTM
The clause weights in Eq. (3) obtain their values during training and can have both positive and negative polarity. It should be noted that no multiplications are needed to implement Eq. (3), only additions, as a clause can output only 0 or 1. The predicted class, $\hat{y}$, is determined by the argmax operator, as given by Eq. (4).
@@ III-C CoTM and Convolution
One can improve the image classification test accuracy by applying convolution [21] to the CoTM. In addition, the number of features that need to be processed simultaneously, is lowered with convolution, which again reduces the model size and the complexity of a design. However, the benefits of convolution come at the expense of increased processing time, which is due to the need to sequentially process the different patches of the input sample.
For image classification, the input data samples to a ConvCoTM, are a set of images, each of dimensions $X\times Y$, with $Z$ channels. The different patches of the image are generated by applying a sliding window of size $W_{X}\times W_{Y}$, with $Z$ channels. The convolution window will be evaluated $B=B_{X}\times B_{Y}$ times across the image, where $B_{X}=1+(X-W_{X})/d_{x}$, $B_{Y}=1+(Y-W_{Y})/d_{y}$, and $d_{x}$ and $d_{y}$ are the stride values of the convolution window in the $X$ and $Y$ directions respectively [21]. For each window position, the number of Boolean features is given by
@@ III-D Datasets and Booleanization
The pixel values range from 0 to 255. For MNIST, they are converted into Boolean variables (i.e., booleanized) with simple thresholding, where pixel values larger than 75 are replaced with 1, and 0 otherwise [21]. For FMNIST and KMNIST, the booleanization is performed with an adaptive Gaussian thresholding procedure [21]. Fig. 1 shows one example of an original image and the corresponding booleanized image from each of the three datasets. For the booleanized images the color white represents a logical 1.
Figure 1: Examples of MNIST, FMNIST and KMNIST image booleanization.
For the booleanization, we use one bit per pixel, i.e., $U=1$ for all three datasets. The convolution window applied has dimensions $W_{X}=W_{Y}=10$, and a step size of 1 is used in both directions. The number of feature bits per patch is therefore 100 bits from the convolution window plus 36 bits for the appended patch position information, ref. Eq. (5). The number of literals for each patch is 272, i.e., two times the number of features.
@@ III-E TM and Convolution - Inference
During inference, each clause in a ConvCoTM is evaluated $B$ times per image. After finishing the convolution, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [21]. The clause output is given by the sequential OR-function in Eq. (6), where $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the window sliding (patch generation).
$c_{j}=\bigvee_{b=0}^{B-1}c_{j}^{b},$ (6)

### v2

# An All-digital 8.6-nJ/Frame 65-nm Tsetlin Machine Image Classification Accelerator

@@ Abstract
We present an all-digital programmable machine learning accelerator chip for image classification, underpinning on the Tsetlin machine (TM) principles. The TM is an emerging machine learning algorithm founded on propositional logic, utilizing sub-pattern recognition expressions called clauses. The accelerator implements the coalesced TM version with convolution, and classifies booleanized images of 28$\times$28 pixels with 10 categories. A configuration with 128 clauses is used in a highly parallel architecture. Fast clause evaluation is achieved by keeping all clause weights and Tsetlin automata (TA) action signals in registers. The chip is implemented in a 65 nm low-leakage CMOS technology, and occupies an active area of 2.7 mm2. At a clock frequency of 27.8 MHz, the accelerator achieves 60.3k classifications per second, and consumes 8.6 nJ per classification. This demonstrates the energy-efficiency of the TM, which was the main motivation for developing this chip. The latency for classifying a single image is 25.4 $\mu$s which includes system timing overhead. The accelerator achieves 97.42%, 84.54% and 82.55% test accuracies for the datasets MNIST, Fashion-MNIST and Kuzushiji-MNIST, respectively, matching the TM software models.
@@ I Introduction
Hardware (HW) accelerators are commonly integrated in modern electronic systems to relieve the burden on system processors from extensive application-specific workloads. Performance and energy-efficiency within a smaller form factor are key considerations for accelerator designs, particularly for Internet-of-Things (IoT), where the compute and memory resources of edge nodes typically are limited. In addition, such nodes are often battery-operated, and it is critical to minimize the overall energy footprints for longer battery operational lifetime.
@@
Current methods for accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3]. CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources. As such, they require large number of micro-architectural operations for each classification and have significant energy footprints [3]. The complexity and energy consumption can be reduced by compromising the data precision [4]. Binary and ternary activations and weights have been applied in several integrated circuit (IC) solutions that have shown good test accuracy with very low power consumption [5, 6, 7, 8, 9].
@@
Unlike the numerical models based on neural networks mentioned above, the Tsetlin machine [10] is an ML algorithm founded on propositional logic. Parallel conjunctive clauses are defined by independent learning automata, and each clause represents a logical sub-pattern. The TM features a single-layer structure with highly interpretable outputs, making it well-suited for application areas such as biomedical, cybernetics, IoT, and law. Furthermore, because its primary operations are of Boolean type, TM solutions are hardware-efficient. The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training. This makes TMs ideal for low-energy application-specific integrated circuit (ASIC) implementations [11]. Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [12].
@@
The original TM version [10] is called the vanilla TM. For image classification, improved test accuracy is achieved by the convolutional TM (CTM) [13]. The CTM achieved a peak test accuracy of 99.4% on the MNIST dataset [14]. For the Fashion-MNIST (FMNIST) [15] and Kuzushiji-MNIST (KMNIST) [16] datasets, the peak test accuracies achieved were 91.5% and 96.31% respectively. Ensemble TM approaches have shown promising accuracies for more complex image datasets, such as CIFAR-10 and CIFAR-100, which are currently being investigated. In the case of CIFAR-10, the approach denoted TM Composites [17] has demonstrated a test accuracy of 82.8% [18].
@@
In this paper, we present an inference accelerator ASIC based on the coalesced TM (CoTM) [19]. The CoTM utilizes a single clause pool, in contrast to the vanilla TM which uses one clause pool per class. As we operate the CoTM with convolution, we here denote it as a convolutional CoTM (ConvCoTM). For configurations with few clauses, the ConvCoTM achieves better accuracy than the vanilla CTM, making it attractive for resource-constrained solutions.
Until now, only a single ASIC based on the TM has been reported [11], which operated on a simple dataset. We wanted to make an an accelerator that operated on a well-known dataset (MNIST) and to demonstrate that an all-digital TM-based ASIC could achieve energy-efficiency at the same level as analog in-memory-computing (IMC) solutions [20, 21, 9].
The ASIC’s accelerator core is built on a modified version of the ConvCoTM inference core included in the FPGA solution in [12], which supported full on-device training. However, because of the FPGA it could not achieve ultra-low-power (ULP) operation. Compared to [12], several enhancements are made in the inference core of the ASIC, as described in section IV. In particular, there is an updated clause logic circuit with a novel feedback that reduces switching activity. There is also new functionality for loading a pre-trained model. Most important is the optimization of the inference part to achieve ULP operation.
Although MNIST is a simple dataset, it is still widely used for benchmarking of ULP image classification ASICs [20, 21, 9]. We therefore consider the MNIST dataset suitable for demonstrating the HW-friendliness and power-efficiency of a ConvCoTM-based accelerator, which was the main motivation for the design and implementation of this ASIC. The main contributions of our paper are:
@@
  • The accelerator is a fully digital, synchronous design, compatible with standard digital workflows. It operates with an energy per classification (EPC) of 8.6 nJ, which is the lowest reported for a fully digital solution running on the same dataset and with comparable test accuracy. Overall, it ranks as the second most energy-efficient solution. Other comparable solutions are of analog and mixed-signal type, utilizing IMC [20, 21, 9].
@@
  • For TMs operating with convolution, we introduce a novel clause logic circuit with feedback, which reduces switching of the clause’s combinational output.
  • We describe an envisaged scaled-up ASIC that operates on the CIFAR-10 dataset, and provide estimates of its performance.
@@ II Related Work
IMC solutions can provide very high energy-efficiency, as the bottleneck related to memory access is relieved. IMC has especially been utilized for analog/mixed-signal ML accelerator designs. An IMC mixed-signal 28 nm CMOS BNN processor is described in [5]. This accelerator achieves an EPC of 3.8 $\mu$J when operating on CIFAR-10. The classification rate is 237 frames per second (FPS) with a power consumption of 0.9 mW.
@@
In [9] a charge-domain IMC ternary neural network (TNN) chip is described. The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model. The accelerator classifies MNIST images at a rate of 549 FPS, consuming 96 $\mu$W. This corresponds to an EPC of 0.18 $\mu$J. The test accuracy achieved is 97.1%.
The accelerator chip that achieves the lowest reported EPC for MNIST classification, is reported in [20]. It is a CNN utilizing an analog IMC architecture with time domain signal processing. It is implemented in 28nm CMOS and achieves an EPC of only 3.32 nJ.
@@
Neuromorphic approaches, utilizing spiking neural networks (SNNs) have achieved very low energy consumption. In [21] a mixed-signal neuromorphic SNN IC is reported, which has been manufactured in a 65 nm CMOS technology. It achieves a test accuracy on MNIST of 95.35%, with a power consumption of 0.517 mW and an EPC of 12.92 nJ. Simulation results of a 28 nm CMOS charge-domain computing SNN solution are reported in [24]. The accelerator achieves an EPC of 15.09 nJ, a latency of 0.46 $\mu$s and a test accuracy on MNIST of 94.81%. In [25], simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described. On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC.
@@
For TM-based HW solutions, the first reported ASIC is described in [11]. It utilizes the vanilla TM architecture and operates on a 3-class machine learning (ML) problem. The chip is implemented in a 65 nm CMOS technology and supports both training and inference. In [26], the first TM-based HW accelerator with convolution is reported. It is an FPGA solution which operates on a 2-class pattern recognition problem (two-dimensional noisy XOR [13]) in 4$\times$4 Boolean images, and employs a 2$\times$2 convolution window. Full on-device training is implemented.
The FPGA solution in [27] is based on the vanilla TM and utilizes the sparsity of typical TM models, in combination with sequential operation. It represents a highly compact solution when it comes to HW resource usage.
@@
In [12] the first reported ConvCoTM-based HW solution is presented, of which the inference part is the basis for our ASIC as described in the Introduction section. It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes full on-device training. For the MNIST dataset a test accuracy of 97.6% is achieved, with a classification rate of 134k samples per second, and an EPC of 13.3 $\mu$J. The source code for this design is available at [28].
@@
In [29, 30, 31] the design and simulation of low-latency asynchronous TM solutions for inference and training are presented. A mixed-signal IMC TM solution is described in [32] with a simulated EPC for MNIST of 13.9 nJ. It is an inference solution with a Boolean-to-current architecture, and resistive RAM (ReRAM) transistor cells are employed to eliminate the need for energy-hungry analog-to-digital and digital-to-analog conversions. Another IMC TM design concept based on Y-flash cells, is reported in [33].
@@ III-A General TM Concepts
A TA of two-action type [34] will either include or exclude a literal in a conjunctive clause, depending on its state. One team of TAs is applied per clause, and the different TA actions are obtained during the training process, which involves feedback mechanisms to each TA [10]. For an inference-only solution, as reported in this paper, only the TA action signals from the trained model are needed, not the complete TAs.
In a TM, the number of clauses, $n$, is a user specified integer. The output of a single clause, $c_{j}$, where $j\in\{0,\ldots,n-1\}$, is given by
@@
where $l_{k}$ is the literal with index k. $k$ belongs to $I_{j}\subseteq\{0,\dots,2o-1\}$, where $I_{j}$ is the set of indices of the literals in clause $j$ for which the corresponding TAs take the action include.
@@ III-B CoTM
The clause weights in Eq. (3) can have both positive and negative polarity, and obtain their values during training. No multiplications are needed to implement Eq. (3), only additions, as a clause can output either 0 or 1. The predicted class, $\hat{y}$, is determined by the argmax operator, as given by Eq. (4).
@@ III-C CoTM and Convolution
It has been shown that applying convolution to the TM can improve the image classification test accuracy [13]. Convolution also reduces the number of features that need to be processed simultaneously, which in turn reduces model size and design complexity. However, the benefits of convolution come at the cost of increased processing time, due to the need to sequentially process the various patches of the input sample.
For image classification, the input data samples to a ConvCoTM, are a set of images, each of dimensions $X\times Y$, with $Z$ channels. The different patches of the image are generated by applying a sliding window of size $W_{X}\times W_{Y}$, with $Z$ channels. The convolution window will be evaluated $B=B_{X}\times B_{Y}$ times across the image, where $B_{X}=1+(X-W_{X})/d_{x}$, $B_{Y}=1+(Y-W_{Y})/d_{y}$, and $d_{x}$ and $d_{y}$ are the stride values of the convolution window in the $X$ and $Y$ directions respectively [13]. For each window position, the number of Boolean features generated per patch is given by
@@
Figure 1: ConvCoTM accelerator ASIC block diagram.
@@ III-D Datasets and Booleanization
The pixel values range from 0 to 255. For MNIST, they are converted into Boolean variables (i.e., booleanized) with simple thresholding, where pixel values larger than 75 are replaced with 1, and 0 otherwise [13]. For FMNIST and KMNIST, an adaptive Gaussian thresholding procedure [13] is applied. We use one bit per pixel, i.e., $U=1$ for all three datasets. The convolution window has dimensions $W_{X}=W_{Y}=10$, and a step size of one is used in both directions. For each patch, we have 100 feature bits from the convolution window plus 36 bits for the appended patch position information, ref. Eq. (5). The number of literals for each patch is therefore 272.
@@ III-E TM and Convolution - Inference
During inference, each clause in a ConvCoTM is evaluated $B$ times per image. After finishing the patch generation, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [13]. The clause output is given by Eq. (6) which we call the sequential OR-function. Here $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the patch generation:
$c_{j}=\bigvee_{b=0}^{B-1}c_{j}^{b}.$ (6)